AO Com Work: 28 October 2010 UT

Observer: DMiller SEsposito, MXompero, LBusoni, APuglisi, FQuiros
Telescope Operator: SAllanson (LBTO)
Telescope Support: MMidkiff (LBTO)
Instrument Support: (LBTO)

Summary

Preparations

Details

00:23 Opened shutters. The AO guys test fit the neutral density filter in front of the pyramid. They did not leave it there, but will be ready to install it later when needed

01:45 pointto M5 star to set IE=-29 CA=29

01:50 pointto AO738 ACTIVE Rmag=7.1

02:00 Collimated to < 400 nm rms Seeing on guider ~ 0.9, DIMM ~ 1.1

02:10 AO loop closed

04:50 iif_preset to AO960 failed for unknown reason (catalog_check error). Had to build target.dat by hand and send it

04:50 Failed preset because no guide star

05:00 Steve pointto Pointing star and set IE=-45 CA=19

05:03 iif_preset AO960 Success Collimated in two interations

05:05 AO closed loop

-- DougMiller - 27 Oct 2010
Topic revision: r3 - 28 Oct 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback